OpenStudy (bookworm14):

Simplify This expression: (I will have to type it in below in the comments sorry)

OpenStudy (anonymous):

lets see book worm!

OpenStudy (bookworm14):

\[\frac{ 4sg }{ -5g }\] A. \[\frac{ -4 }{ 5 }s\] B. \[\frac{ 4 }{ 5 }g\] C. \[\frac{ -5 }{ 4 }s\] D. \[\frac{ -5 }{ 4 }g\] PLEASE HELP !!!

OpenStudy (anonymous):

what math is this?

OpenStudy (bookworm14):

Algebra 1, why do you ask?

OpenStudy (bookworm14):

Ive done these before and passed it, but I kinda forgot how to do it

OpenStudy (anonymous):

pre-calc student and just curious

OpenStudy (anonymous):

tbh i dont really remember how to do this either, but my guess would be C.) -5/4s

OpenStudy (anonymous):

my best guess*

OpenStudy (cj49):

do u still need help..

OpenStudy (cj49):

The g/g cancels out the g's. 4s/-5 - 4s/5

OpenStudy (bookworm14):

that was the correct answer @cj49 thank you for helpin me !!

OpenStudy (cj49):

np c: